### v1

@@ Abstract
We propose the Regge trajectory relations for the fully heavy pentaquark $P_{cc\bar{c}bb}$ utilizing both diquark and triquark Regge trajectory relations. Using these new relations, we discuss four series of Regge trajectories: the $\rho_{1}$-, $\rho_{2}$-, $\lambda_{1}$-, and $\lambda_{2}$-trajectories. We provide rough estimates for the masses of the $\rho_{1}$-, $\rho_{2}$-, $\lambda_{1}$-, and $\lambda_{2}$-excited states. Except for the $\lambda_{1}$-trajectories, the complete forms of the other three series of Regge trajectories for the pentaquark $P_{cc\bar{c}bb}$ are lengthy and cumbersome. We show that the $\rho_{1}$-, $\rho_{2}$-, and $\lambda_{2}$-trajectories can not be obtained by simply imitating the meson Regge trajectories because mesons have no substructures. To derive these trajectories, pentaquark’s structure and substructure should be taken into consideration. Otherwise, the $\rho_{1}$-, $\rho_{2}$-, and $\lambda_{2}$-trajectories must rely solely on fitting existing theoretical or future experimental data. The fundamental relationship between the slopes of the obtained trajectories and constituents’ masses and string tension will remain unclear, and the predictive power of the Regge trajectories would be compromised. Moreover, we show that the lengthy complete forms of the $\rho_{1}$-, $\rho_{2}$-, and $\lambda_{2}$-trajectories can be well approximated by the simple fitted formulas. Four series of Regge trajectories for the pentaquark $P_{cc\bar{c}bb}$ all exhibit a behavior of $M{\sim}x^{2/3}$, where $x=n_{r_{1}},n_{r_{2}},l_{1},l_{2},N_{r_{1}},N_{r_{2}},L_{1},L_{2}$. All four series of trajectories exhibit concave downward behavior in the $(M^{2},\,x)$ plane.
@@ I Introduction
The Regge trajectory [^1] A Regge trajectory of bound states is generally expressed as $M=m_{R}+\beta_{x}(x+c_{0})^{\nu}$ $(x=l,\,n_{r})$ Chen:2022flh ; Chen:2021kfw , where $M$ is the mass of the bound state, $l$ is the orbital angular momentum, and $n_{r}$ is the radial quantum number. $m_{R}$ and $\beta_{x}$ are parameters. For simplicity, plots in the $(M,\,x)$ plane Xie:2024lfo , $(M-m_{R},\,x)$ plane Chen:2023cws , $(M,\,(x+c_{0})^{\nu})$ plane Burns:2010qq , $(M^{2},\,x)$ plane Chen:2018nnr , $((M-m_{R})^{2},\,x)$ plane Chen:2023djq ; Chen:2023web or $((M-m_{R})^{1/{\nu}},\,x)$ plane Xie:2024dfe , are all commonly referred as Chew-Frautschi plots. Regge trajectories can be plotted in these various planes. is one of the effective approaches widely used in the study of hadron spectra Burns:2010qq ; Regge:1959mz ; Chew:1962eu ; Nambu:1978bd ; Gross:2022hyw ; Brodsky:2006uq ; Nielsen:2018uyn ; brau:04bs ; Brisudova:1999ut ; Guo:2008he ; Ebert:2009ub ; Irving:1977ea ; Collins:1971ff ; Inopin:1999nf ; Afonin:2014nya ; MartinContreras:2020cyg ; Sergeenko:1994ck ; Veseli:1996gy ; Inopin:2001ub ; Wilczek:2004im ; Sonnenschein:2018fph ; MartinContreras:2023oqs ; Roper:2024ovj ; G:2024zkc ; Oudichhya:2024ikt . Few studies have applied the Regge trajectory approach to the discussion of pentaquarks. In Ref. Ghosh:2017cck , Regge trajectories for the pentaquarks containing one heavy quark were investigated. In Ref. Sindhu:2023oqo , Regge trajectories for the pentaquarks containing two charm quarks were discussed. In both of these references, only one series of Regge trajectories are discussed. To our knowledge, there have been no systematic studies up to now addressing both two series of $\lambda$-trajectories and two series of $\rho$-trajectories for pentaquarks. In the diquark-triquark picture, a pentaquark consists of a diquark and a triquark, where the triquark is composed of an antiquark and a diquark (see Fig. 1). By employing the diquark Regge trajectory relations Feng:2023txx and the triquark Regge trajectory relations Song:2024bkj , we attempt to apply the Regge trajectory approach to the pentaquark $P_{cc\bar{c}bb}$. The masses of two $\lambda$-mode and two $\rho$-mode excited states are roughly estimated. Four series of Regge trajectories–the $\lambda_{1}$-, $\lambda_{2}$-, $\rho_{1}$-, and $\rho_{2}$-trajectories–are investigated.
The paper is organized as follows: In Sec. II, the Regge trajectory relations for the the fully heavy pentaquarks are proposed. In Sec. III, four series of masses and four series of Regge trajectories are presented. In Sec. IV, the discussions are given. The conclusions are presented in Sec. V.
@@ II.1 Preliminary
Figure 1: Schematic diagram of a pentaquark in the diquark-triquark picture. The left grey part represents the diquark $1$, composed of two quarks ($q_{1}$ and $q_{2}$). The right grey part represents a triquark composed of one antiquark ($\bar{q}_{3}$) and diquark $2$. Diquark $2$ is composed of quarks $q_{4}$ and $q_{5}$. The circles denote the quarks and the antiquark.
@@
In $SU_{c}(3)$, $3_{c}\otimes 3_{c}=\bar{3}_{c}\oplus{6}_{c}$. There is attraction between quark pairs $(q_{1}q_{2})$ in the color antitriplet channel. Only the $\bar{3}_{c}$ representation of $SU_{c}(3)$ is considered in the present work; the $6_{c}$ representation is not considered. Similarly, in the diquark picture, a triquark $(\bar{q}_{3}(q_{4}q_{5}))$ is regarded as a bound state consisting of a diquark $(q_{4}q_{5})$ in $\bar{3}_{c}$ and an antiquark $\bar{q}_{3}$ also in $\bar{3}_{c}$ Lebed:2015tna . In the triquark-diquark model, the pentaquarks under consideration are composed of a diquark and a triquark in the color $\bar{3}_{c}$ and $3_{c}$ configurations.
@@
[…]
@@ II.2 Regge trajectory relations for the fully heavy pentaquarks
$M\Psi_{d,t,p}({{\bf r}})=\left(\omega_{1}+\omega_{2}\right)\Psi_{d,t,p}({{\bf r}})+V_{d,t,p}\Psi_{d,t,p}({{\bf r}}),$ (2)
@@
The effect of the finite size of the diquark is treated differently. In Ref. Faustov:2021hjs , the size of diquark is taken into account through appropriate form factors, while sometimes the diquark is treated as pointlike Ferretti:2019zyh ; Lundhammar:2020xvw . In the present work, the spinless Salpeter equation (2) is not used to calculate the masses of bound states but rather to obtain the Regge trajectory relations. For simplicity, the diquark is regarded as pointlike.
@@
where $V_{c}\propto{1/r}$ is a color Coulomb potential or a Coulomb-like potential arising from one-gluon-exchange. Here, $\sigma$ is the string tension, and $C$ is a fundamental parameter Gromes:1981cb ; Lucha:1991vn . The expression in brackets is the Cornell potential Eichten:1974af . The first two terms in Eq. (4) are determined solely by the nature of the color field; therefore, the same potential used for mesons and baryons Godfrey:1985xj ; Capstick:1986ter can also be used to describe the interaction between diquark and triquark Lebed:2015tna ; Brodsky:2014xia ; Song:2024bkj . ${\bf{F}_{i}}\cdot{\bf{F}_{j}}$ is the color-Casimir, with
$\langle{(\bf{F}_{i}}\cdot{\bf{F}_{j}})_{d,t}\rangle=-\frac{2}{3},\quad\langle{(\bf{F}_{i}}\cdot{\bf{F}_{j}})_{p}\rangle=-\frac{4}{3}.$ (5)
@@
In Eq. (II.2), $M_{d_{1}}$ and $M_{d_{2}}$ give the diquark Regge trajectories Feng:2023txx , $M_{d_{1}}$, $M_{d_{2}}$ and $M_{t}$ can can be used to construct the triquark Regge trajectories Song:2024bkj . $M_{d_{1}}$, $M_{d_{2}}$, $M_{t}$ and $M$ generate the Regge trajectories for the fully heavy pentaquarks. According to Eqs. (II.2) and (II.2), we have
@@ III.2 $\rho$-trajectories for the pentaquark $P_{cc\bar{c}bb}$
Table 2: The masses of the radially and orbitally $\rho_{1}$- and $\rho_{2}$-excited states of the pentaquark $P_{cc\bar{c}bb}$ (in ${{\rm GeV}}$). Both configurations $(cc)(\bar{c}(bb))$ and $(bb)(\bar{c}(cc))$ are considered. The notation in Eq. (1) is rewritten as $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N_{2}^{2s_{3}+1}L_{2J_{2}},N_{1}^{2s_{4}+1}L_{1J_{1}}\rangle$. $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N_{2}L_{2},N_{1}L_{1}\rangle$ denotes the spin-averaged states. Eqs. (II.2), (II.2) and (17) are used in the calculation. States marked with (×)($\times$)( × ) do not exist.
$|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N_{2}L_{2},N_{1}L_{1}\rangle$ $(cc)(\bar{c}(bb))$ $(bb)(\bar{c}(cc))$
@@
where $M\left(\rho_{1},(bb)(\bar{c}(cc))\right)$ denotes the masses of the radially or orbitally $\rho_{1}$-mode excited state of $P_{cc\bar{c}bb}$ in the configuration $(bb)(\bar{c}(cc))$; the other notations in Eq. (III.2) are similarly defined. The masses of the $\rho_{1}$-excited ($\rho_{2}$-excited) states in the configuration $(cc)(\bar{c}(bb))$ are close to those of the $\rho_{2}$-excited ($\rho_{1}$-excited) states in the configuration $(bb)(\bar{c}(cc))$, i.e., $M\big{(}\rho_{1},(cc)(\bar{c}(bb))\big{)}{\approx}M\big{(}\rho_{2},(bb)(\bar{c}(cc))\big{)}$, $M\big{(}\rho_{1},(bb)(\bar{c}(cc))\big{)}{\approx}M\big{(}\rho_{2},(cc)(\bar{c}(bb))\big{)}$. Possible mixings between these states are neglected here.
@@
Inspired by Ref. Burns:2010qq , we plot the $\rho_{1}$- and $\rho_{2}$-trajectories for the pentaquark $P_{cc\bar{c}bb}$ in configurations $(cc)(\bar{c}(bb))$ and $(bb)(\bar{c}(cc))$ in the $(M,\,(x+c_{0})^{2/3})$ plane (see Fig. 2). The full forms of the Regge trajectories are directly calculated from Eqs. (II.2) and (II.2), or from Eqs. (II.2) and (II.2). As examples, the full forms of the $\rho_{1}$-trajectories and $\rho_{2}$-trajectories are listed in Eqs. (A) and (A). These expressions are generally lengthy and cumbersome. By linearly fitting the calculated data in Table 2, we obtain the fitted formulas listed in Table 5. As shown in Fig. 2, the fitted formulas overlap with the complete forms of the $\rho_{1}$- and $\rho_{2}$-trajectories. This indicates that the complicated full forms of the $\rho$-Regge trajectories can be well approximated by the simple fitted formulas. Therefore, we have that the complete forms of the $\rho_{1}$- and $\rho_{2}$-trajectories all exhibit the behavior $M{\sim}x_{\rho_{1}}^{2/3},\;x_{\rho_{2}}^{2/3}$, where $x_{\rho_{1}}=n_{r_{1}},\;l_{1}$ and $x_{\rho_{2}}=n_{r_{2}},\;l_{2}$. The main parts of the full forms of the $\rho_{1}$- and $\rho_{2}$-trajectories [listed in Table 5] show significant deviations from the full forms. For both the radial and orbital Regge trajectories, and for both configurations $(cc)(\bar{c}(bb))$ and $(bb)(\bar{c}(cc))$, the main parts of the $\rho_{2}$-trajectories deviate more from the complete forms than those of the $\rho_{1}$-trajectories.

### v3

@@ Abstract
We propose the Regge trajectory relations for the fully heavy pentaquark $P_{cc\bar{c}bb}$ utilizing both diquark and triquark Regge trajectory relations. Using these new relations, we discuss four series of Regge trajectories: the $\rho_{1}$-, $\rho_{2}$-, $\lambda_{1}$-, and $\lambda_{2}$-trajectories. We provide rough estimates for the masses of the $\rho_{1}$-, $\rho_{2}$-, $\lambda_{1}$-, and $\lambda_{2}$-excited states. Except for the $\lambda_{1}$-trajectories, the complete forms of the other three series of Regge trajectories for the pentaquark $P_{cc\bar{c}bb}$ are lengthy and cumbersome. We show that the $\rho_{1}$-, $\rho_{2}$-, and $\lambda_{2}$-trajectories can not be obtained by simply imitating the meson Regge trajectories because mesons have no substructures. To derive these trajectories, pentaquark’s structure and substructure should be taken into consideration. Otherwise, the $\rho_{1}$-, $\rho_{2}$-, and $\lambda_{2}$-trajectories must rely solely on fitting existing theoretical or future experimental data. Consequently, the fundamental relationship between the slopes of the obtained trajectories and constituents’ masses and string tension will become unobvious, and the predictive power of the Regge trajectories would be compromised. Moreover, we show that the lengthy complete forms of the $\rho_{1}$-, $\rho_{2}$-, and $\lambda_{2}$-trajectories can be well approximated by the simple fitted formulas. Four series of Regge trajectories for the pentaquark $P_{cc\bar{c}bb}$ all exhibit a behavior of $M{\sim}x^{2/3}$, where $x=n_{r_{1}},n_{r_{2}},l_{1},l_{2},N_{r_{1}},N_{r_{2}},L_{1},L_{2}$. All four series of trajectories exhibit concave downward behavior in the $(M^{2},\,x)$ plane.
@@ I Introduction
The Regge trajectory [^1] A Regge trajectory of bound states is generally expressed as $M=m_{R}+\beta_{x}(x+c_{0})^{\nu}$ $(x=l,\,n_{r})$ Chen:2022flh ; Chen:2021kfw , where $M$ is the mass of the bound state, $l$ is the orbital angular momentum, and $n_{r}$ is the radial quantum number. $m_{R}$ and $\beta_{x}$ are parameters. For simplicity, plots in the (M,x)(M,\,x) plane Xie:2024lfo , $(M-m_{R},\,x)$ plane Chen:2023cws , $(M,\,(x+c_{0})^{\nu})$ plane Burns:2010qq , $(M^{2},\,x)$ plane Chen:2018nnr , $((M-m_{R})^{2},\,x)$ plane Chen:2023djq ; Chen:2023web or $((M-m_{R})^{1/{\nu}},\,x)$ plane Xie:2024dfe , are all commonly referred as Chew-Frautschi plots. Regge trajectories can be plotted in these various planes. is one of the effective approaches widely used in the study of hadron spectra Burns:2010qq ; Regge:1959mz ; Chew:1962eu ; Nambu:1978bd ; Gross:2022hyw ; Brodsky:2006uq ; Nielsen:2018uyn ; brau:04bs ; Brisudova:1999ut ; Guo:2008he ; Ebert:2009ub ; Irving:1977ea ; Collins:1971ff ; Inopin:1999nf ; Afonin:2014nya ; MartinContreras:2020cyg ; Sergeenko:1994ck ; Veseli:1996gy ; Inopin:2001ub ; Wilczek:2004im ; Sonnenschein:2018fph ; MartinContreras:2023oqs ; Roper:2024ovj ; G:2024zkc ; Oudichhya:2024ikt . Few studies have applied the Regge trajectory approach to the discussion of pentaquarks. In Ref. Ghosh:2017cck , Regge trajectories for the pentaquarks containing one heavy quark were investigated. In Ref. Sindhu:2023oqo , Regge trajectories for the pentaquarks containing two charm quarks were discussed. In both of these references, only one series of Regge trajectories are discussed. To our knowledge, there have been no systematic studies up to now addressing both two series of $\lambda$-trajectories and two series of $\rho$-trajectories for pentaquarks. In the diquark-triquark picture, a pentaquark consists of a diquark and a triquark, where the triquark is composed of an antiquark and a diquark (see Fig. 1). By employing the diquark Regge trajectory relations Feng:2023txx and the triquark Regge trajectory relations Song:2024bkj , we attempt to apply the Regge trajectory approach to the pentaquark $P_{cc\bar{c}bb}$. The masses of two $\lambda$-mode and two $\rho$-mode excited states are roughly estimated. Four series of Regge trajectories–the $\lambda_{1}$-, $\lambda_{2}$-, $\rho_{1}$-, and $\rho_{2}$-trajectories–are investigated.
The paper is organized as follows: In Sec. II, the Regge trajectory relations for the fully heavy pentaquarks are proposed. In Sec. III, four series of masses and four series of Regge trajectories are presented. In Sec. IV, the discussions are given. The conclusions are presented in Sec. V.
@@ II.1 Preliminary
Figure 1: Schematic diagram of a pentaquark in the diquark-triquark picture. The left grey part represents the diquark $1$, composed of two quarks ($q_{1}$ and $q_{2}$). The right grey part represents a triquark composed of one antiquark ($\bar{q}_{3}$) and diquark $2$. Diquark 22 is composed of quarks $q_{4}$ and $q_{5}$. The circles denote the quarks and the antiquark.
@@
A diquark $(q_{1}q_{2})$ can couple to only two irreducible color representations: $3_{c}\otimes 3_{c}=\bar{3}_{c}\oplus{6}_{c}$. The $\bar{3}_{c}$ is the attractive channel, while in the $6_{c}$ representation, the internal interaction between the $q_{1}q_{2}$ pair is repulsive. In some works, both $\bar{3}_{c}$ and ${6}_{c}$ are considered, for example, see Refs. Maiani:2019lpu ; Berwein:2024ztx , while in other works, only $\bar{3}_{c}$ is considered, for example, see Refs. Brodsky:2014xia ; Galkin:2023wox . Following Refs. Brodsky:2014xia ; Galkin:2023wox , only the $\bar{3}_{c}$ diquark is considered in the present work. Similarly, the triquark under consideration, $(\bar{q}_{3}(q_{4}q_{5}))$, is a ${3}_{c}$ bound state consisting of a diquark $(q_{4}q_{5})$ in the $\bar{3}_{c}$ representation and an antiquark $\bar{q}_{3}$ also in the $\bar{3}_{c}$ representation. In the triquark-diquark model, the color-singlet pentaquarks under consideration are composed of a diquark in $\bar{3}_{c}$ and a triquark in $3_{c}$ Lebed:2015tna .
@@
where $\bar{3}_{c}$ denotes the color antitriplet state of the diquark, and $1_{c}$ represents the color singlet state of the pentaquark. [The superscript $1_{c}$ is often omitted, as the observed pentaquarks are colorless.] The notation in Eq. (1) can also be written as $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N_{2}^{2s_{3}+1}L_{2J_{2}},N_{1}^{2s_{4}+1}L_{1J_{1}}\rangle$. The diquark $(q_{1}q_{2})$ is either $\{q_{1}q_{2}\}$ or $q_{1}q_{2}]$, where $\{q_{1}q_{2}\}$ and $q_{1}q_{2}]$ represent the permutation symmetric and antisymmetric flavor wave functions, respectively. $N_{1,2}=N_{r_{1,2}}+1$, where $N_{r_{1,2}}=0,\,1,\,\cdots$. $n_{1,2}=n_{r_{1,2}}+1$, where $n_{r_{1,2}}=0,\,1,\,\cdots$. $N_{r_{1}}$, $N_{r_{2}}$, $n_{r_{1}}$ and $n_{r_{2}}$ are the radial quantum numbers of the pentaquark, triquark, diquark 1, diquark 2, respectively. $\vec{J}_{1}=\vec{L}_{1}+\vec{s}_{4}$, $\vec{s}_{4}=\vec{j}_{1}+\vec{J}_{2}$, $\vec{J}_{2}=\vec{s}_{3}+\vec{L}_{2}$, $\vec{s}_{3}=\vec{j}_{2}+\vec{s}_{\bar{q}_{3}}$, $\vec{j}_{1}=\vec{s}_{1}+\vec{l}_{1}$, $\vec{j}_{2}=\vec{s}_{2}+\vec{l}_{2}$. $\vec{J}_{1}$, $\vec{J}_{2}$, $\vec{j}_{1}$ and $\vec{j}_{2}$ are the spins of pentaquark, triquark, diquark 1, diquark 2, respectively. $L_{1}$, $L_{2}$, $l_{1}$ and $l_{2}$ are the orbital quantum numbers of pentaquark, triquark, diquark 1 and diquark 2, respectively. $\vec{s}_{1}$ and $\vec{s}_{2}$ are the summed spin of quarks in the diquark $1$ and diquark $2$, respectively; $\vec{s}_{3}$ is the summed spin of antiquark $3$ and diquark $2$, and $\vec{s}_{4}$ is the summed spin of diquark $1$ and triquark.
@@ II.2 Regge trajectory relations for the fully heavy pentaquarks
$M\Psi_{d,t,p}({{\bf r}})=H\Psi_{d,t,p}({{\bf r}}),\;\,H=\omega_{1}+\omega_{2}+V_{d,t,p},$ (2)
@@
The effect of the finite size of the diquark is treated differently. In Ref. Faustov:2021hjs , the size of diquark is taken into account through appropriate form factors, while sometimes the diquark is treated as pointlike Ferretti:2019zyh ; Lundhammar:2020xvw . For simplicity, the diquark is regarded as pointlike in the present work.
@@
where $V_{c}\propto{1/r}$ is a color Coulomb potential or a Coulomb-like potential arising from one-gluon-exchange. $C$ is a fundamental parameter Gromes:1981cb ; Lucha:1991vn . $\sigma$ is the string tension. The string tension $\sigma$ can be associated with the constituent quark masses Burns:2010qq ; Lucha:1991vn or treated as a common parameter for different quark masses Faustov:2021hjs ; Faustov:2021qqf ; Lucha:2000hg . In this work, we adopt the latter approach. Although the masses for $J/\Psi$ and $\Psi^{\prime}$ have undergone significant improvements since the 1970s, we use the value $\sigma=0.18$ ${\rm GeV}^{2}$ Godfrey:1985xj ; Ebert:2007rn ; Faustov:2021hjs . ${\bf{F}_{i}}\cdot{\bf{F}_{j}}$ is the color-Casimir, with
$\langle{(\bf{F}_{i}}\cdot{\bf{F}_{j}})_{d,t}\rangle=-\frac{2}{3},\quad\langle{(\bf{F}_{i}}\cdot{\bf{F}_{j}})_{p}\rangle=-\frac{4}{3},$ (5)
where $\langle{(\bf{F}_{i}}\cdot{\bf{F}_{j}})_{d,t,p}\rangle$ are the color-Casimirs for the diquark, the triquark, and the pentaquark, respectively.
The expression in brackets in Eq. (4) is the Cornell potential. The dynamics of hadrons are governed by a Hamiltonian that features one-gluon exchange dominance at short distances and a universal confinement potential independent of the hadron’s quark content and dominating at large distances. This framework can accurately describe the spectroscopy of heavy mesons and baryons Ferretti:2019zyh ; Godfrey:1985xj ; Eichten:1974af . The Cornell form also persists in multiquark systems, including tetraquarks in the diquark-antidiquark picture Ferretti:2019zyh ; Bedolla:2019zwg ; Xie:2024dfe and pentaquarks in the diquark-triquark picture Lebed:2015tna . In the diquark picture, a baryon consists of one quark in $3_{c}$ and a diquark in $\bar{3}_{c}$, while a pentaquark is composed of a diquark in $\bar{3}_{c}$ and a triquark in $3_{c}$. The Cornell approach reduces the pentaquark modeling (via diquark-triquark interaction) problem to a meson-like system, where a meson is composed of a quark in $3_{c}$ and an antiquark in $\bar{3}_{c}$ Ferretti:2019zyh ; Bedolla:2019zwg ; Lebed:2015tna ; Brodsky:2014xia ; Song:2024bkj . The color-Casimir values corresponding to quark-antiquark interactions for mesons, quark-diquark interactions for baryons, antidiquark-diquark interactions for tetraquarks, and triquark-diquark interactions for pentaquarks are all 4/3-4/3.
@@
In Eq. (II.2), $M_{d_{1}}$ and $M_{d_{2}}$ give the diquark Regge trajectories Feng:2023txx , $M_{d_{1}}$, $M_{d_{2}}$ and $M_{t}$ can be used to construct the triquark Regge trajectories Song:2024bkj . $M_{d_{1}}$, $M_{d_{2}}$, $M_{t}$ and $M$ generate the Regge trajectories for the fully heavy pentaquarks. According to Eqs. (II.2) and (II.2), we have
@@
In the diquark-triquark picture, the problem of modeling pentaquarks (via diquark-triquark interaction) reduces to a meson-like system Lebed:2015tna ; Brodsky:2014xia . In the meson case, spin-dependent interactions are relativistic corrections and are small compared to the Cornell potential. These contributions do not alter the behavior of meson Regge trajectories. Similarly, spin-dependent interactions between the diquark and triquark do not affect the behavior of pentaquark Regge trajectories. In this work, we focus on the Regge trajectories of fully heavy pentaquarks and do not consider spin-dependent effects, which will be addressed in future studies.
@@ III.2 $\rho$-trajectories for the pentaquark $P_{cc\bar{c}bb}$
Table 2: The masses of the radially and orbitally $\rho_{1}$- and $\rho_{2}$-excited states of the pentaquark $P_{cc\bar{c}bb}$. Both configurations $(cc)(\bar{c}(bb))$ and $(bb)(\bar{c}(cc))$ are considered. The notation in Eq. (1) is rewritten as $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N_{2}^{2s_{3}+1}L_{2J_{2}},N_{1}^{2s_{4}+1}L_{1J_{1}}\rangle$. $|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N_{2}L_{2},N_{1}L_{1}\rangle$ denotes the spin-averaged states. Eqs. (II.2), (II.2) and (17) are used in the calculation. States marked with $(\times)$ do not exist.
$|n_{1}^{2s_{1}+1}l_{1j_{1}},n_{2}^{2s_{2}+1}l_{2j_{2}},N_{2}L_{2},N_{1}L_{1}\rangle$ $(cc)(\bar{c}(bb))$ [GeV] $(bb)(\bar{c}(cc))$ [GeV]
@@
where $M\left(\rho_{1},(bb)(\bar{c}(cc))\right)$ denotes the masses of the radially or orbitally $\rho_{1}$-mode excited state of $P_{cc\bar{c}bb}$ in the configuration $(bb)(\bar{c}(cc))$; the other notations in Eq. (18) are similarly defined. The masses of the $\rho_{1}$-excited ($\rho_{2}$-excited) states in the configuration $(cc)(\bar{c}(bb))$ are close to those of the $\rho_{2}$-excited ($\rho_{1}$-excited) states in the configuration $(bb)(\bar{c}(cc))$, i.e., $M\big(\rho_{1},(cc)(\bar{c}(bb))\big){\approx}M\big(\rho_{2},(bb)(\bar{c}(cc))\big)$, $M\big(\rho_{1},(bb)(\bar{c}(cc))\big){\approx}M\big(\rho_{2},(cc)(\bar{c}(bb))\big)$. Possible mixings between these states are neglected here.
Unlike the Regge trajectories for light mesons, which have a simple linear form, the Regge trajectories for fully heavy pentaquarks become tedious when plotted in the $(M^{2},x)$ plane. For example, the fitted $\rho_{1}$-Regge trajectory changes from Eq. (23) to
$M^{2}=$ $195.052+0.216964(0.205+n_{r_{1}})^{4/3}$
$+13.0107(0.205+n_{r_{1}})^{2/3}.$ (19)
The squared complete form of $\rho_{1}$-Regge trajectory [see Eq. (A)] will be even more complicated. Furthermore, the behavior of the Regge trajectory in terms of $M^{2}$ is less obvious and transparent than in terms of $M$. For the fitted $\rho_{1}$-Regge trajectory, $M{\sim}n_{r_{1}}^{2/3}$, while $M^{2}{\sim}n_{r_{1}}^{\nu}$ with $\nu$ not being clearly defined but greater than $2/3$, and $\nu$ varies for different trajectories. When Eq. (III.2) is approximated as $M^{2}=m^{\prime}_{R}+\beta^{\prime}(x+c_{0})^{2/3}$ because the term $(0.205+n_{r_{1}})^{2/3}$ plays the dominant role, the resulting estimates become less accurate compared with those obtained using $M=m_{R}+\beta(x+c_{0})^{2/3}$.
@@
Inspired by Ref. Burns:2010qq , we employ the $(M,\,(x+c_{0})^{2/3})$ plane, rather than the $(M^{2},\,x)$ plane, to plot the $\rho_{1}$- and $\rho_{2}$-trajectories for the pentaquark $P_{cc\bar{c}bb}$ in the configurations $(cc)(\bar{c}(bb))$ and $(bb)(\bar{c}(cc))$ (see Fig. 2). The full forms of the $\rho_{1}$- and $\rho_{2}$-trajectories, directly calculated from Eqs. (II.2) and (II.2) or from Eqs. (II.2) and (II.2), are indicated by the black dashed lines in Fig. 2. For illustration, the full expressions are presented in Eqs. (A) and (A). These expressions are generally lengthy and cumbersome, and their complexity varies with different Regge trajectories. By performing a linear fit to the calculated data in Table 2, we obtain the fitted formulas listed in Table 6. In Fig. 2, the fitted formulas are represented by the pink lines. As shown in Fig. 2, these fitted formulas (pink lines) nearly overlap with the complete forms of the $\rho_{1}$- and $\rho_{2}$-trajectories (black dashed lines). This demonstrates that the complex full forms of the $\rho$-Regge trajectories can be well approximated by the simple fitted formulas. Accordingly, the complete forms of the $\rho_{1}$- and $\rho_{2}$-trajectories both exhibit the behavior $M{\sim}x_{\rho_{1}}^{2/3},\;x_{\rho_{2}}^{2/3}$, where $x_{\rho_{1}}=n_{r_{1}},\;l_{1}$ and $x_{\rho_{2}}=n_{r_{2}},\;l_{2}$. The main parts of the full forms of the $\rho_{1}$- and $\rho_{2}$-trajectories (listed in Table 6 and indicated by the blue dotted lines in Fig. 2) show significant deviations from the full forms (black dashed lines). For both the radial and orbital Regge trajectories, and for both configurations $(cc)(\bar{c}(bb))$ and $(bb)(\bar{c}(cc))$, the main parts of the $\rho_{2}$-trajectories deviate more from the complete forms than do those of the $\rho_{1}$-trajectories. In conclusion, the $\rho$-trajectories are presented in Fig. 2, and the Chew-Frautschi plots clearly demonstrate the behavior of trajectories (see the footnote on the first page). The tedious complete forms of the $\rho$-trajectories–in which the Regge trajectory behavior is not obvious–can be well approximated by the simple fitted forms, in which the Regge trajectory behavior is apparent. Furthermore, the differences between the main parts and the complete forms vary among different series of Regge trajectories.